OpenStudy (anonymous):

Electronic baseball games manufactured by Tempco Electronics are shipped in lots of 36. Before shipping, a quality-control inspector randomly selects a sample of 7 from each lot for testing. If the sample contains any defective games, the entire lot is rejected. What is the probability that a lot containing exactly 2 defective games will still be shipped? (Round your answer to three decimal places.)

OpenStudy (zarkon):

hypergeometric distribution

OpenStudy (kropot72):

Proportion defective = 2/36 = p P is the chance of acceptance = P \[P=(1-p)^{7}=(1-\frac{2}{36})^{7}=0.67\]

OpenStudy (zarkon):

I assume the sampling is done without replacement. ? ?

OpenStudy (kropot72):

This is single sampling.

OpenStudy (zarkon):

is the sampling done where you pick 7 different items...or is it possible to select the same item more than once.

OpenStudy (kropot72):

In this type of sampling the sample is selected randomly (say using a random number generator). If the same number comes again during the selection process then discard the result and select again

OpenStudy (zarkon):

then the sampling is done without replacement...use the hypergeometric distribution

OpenStudy (kropot72):

I understand your reply based on my answer to your question about the method of sampling. My answer is wrong. To ensure that sampling is truly random a table of random numbers should be used. These consist of digits in an order which has been tested and found to be free from any systematic defects. A number is assigned to each unit liable to be tested. Then reference to the table will ensure a random selection is made. The question states that that the inspector "randomly selects a sample". If truly random sampling was used then I believe my answer to the question is correct.

OpenStudy (zarkon):

your answer is correct if we are randomly sampling with replacement

OpenStudy (kropot72):

The question deals with lot acceptance sampling without replacement. The lot size is small so in this case you are right, the hypergeometric distribution applies.

OpenStudy (kropot72):

I used the hypergeometric calculator at Stat Trek. This calculator gave the hypergeometric probability of finding one defective game as follows: P(x=1) = 0.3222 Therefore the probability that the defective lot will be shipped is: P(acceptance) = 1 - 0.3222 = 0.678

OpenStudy (zarkon):

the probability is 29/45

OpenStudy (kropot72):

Zarkon, Did you calculate using matrices?

OpenStudy (zarkon):

\[\frac{{34\choose 7}{2\choose 0}}{{36\choose 7}}\]

OpenStudy (kropot72):

\[\frac{\left(\begin{matrix}34 \\ 6\end{matrix}\right)\left(\begin{matrix}2 \\ 1\end{matrix}\right)}{\left(\begin{matrix}36\\ 7\end{matrix}\right)}\] I expected this.

OpenStudy (zarkon):

the question is ...What is the probability that a lot containing exactly 2 defective games will still be shipped? in order to be shipped there can't be any defectives...therefore out of the 2 defective we want 0 of them